# Differential Privacy for Transformer Embeddings with Nonparametric Variational Information Bottleneck

### v1

@@ 4.2 Results on the Glue Benchmark
Experimental Protocol.
For each model, we perform five independent runs and select the best-performing run on the validation set for final evaluation on the test set. The NVDP and VIB models provide privacy by applying the same learned stochastic mapping at test time, ensuring that only noisy sanitized embeddings are shared. The utility of these private embeddings is measured by evaluating the classifier learned during training. The results in Table 1 represent the best privacy-utility trade-off found by sweeping across a range of regularization weights $\lambda_{G}$ and $\lambda_{D}$ (as defined in Equation 5). To quantify the privacy loss shown in Table 1, we fix the Rényi order to $\lambda=1.1$ (in Equation 7) and the BDP failure probability to $\delta_{\mu}=10^{-5}$ (in Equation 3) and reporting both the Bayesian Differential Privacy (BDP) and the maximum Rényi Divergence (RD max) across all test set pairs. Full results for all tested regularization strengths are provided in Appendix A.
Table 1: Privacy-utility trade-off on GLUE tasks for BERT-Base. We compare our proposed NVDP model against non-private baselines and private ablations (VIB-fixed, VIB-learned, and VTDP). For each private model, we report its best-achieved utility score alongside privacy guarantees. Privacy is measured via Bayesian Differential Privacy (BDP $\downarrow$) and Rényi Divergence (RD $\downarrow$). Lower privacy values are better. The best-performing private model per task is bolded.
@@
Utility and Regularization Analysis.
@@
We analyze this ability of our models to balance utility against privacy, to get a more complete understanding of the results from Table 1. Figure 2 plots accuracy against the BDP measure for the tunable models (NVDP and VTDP); detailed results are given in Table 2 in Appendix A. For this Bayesian Differential Privacy guarantee, the results show a clear advantage for our proposed model. Across all datasets, the NVDP models (blue curves) consistently occupy the most favorable region of the plot – closest to the top-right corner – indicating a superior trade-off compared to the VTDP ablation (red curves). For example, on MRPC, to achieve a privacy budget comparable to NVDP (e.g., $\approx$ 10.6), the VTDP ablation’s accuracy drops to just 74.8%.
Figure 2: Accuracy versus Bayesian Differential Privacy ($\epsilon_{\mu}$). The BDP budget ($\epsilon_{\mu}$) is calculated by finding the tightest privacy guarantee for a fixed $\delta_{\mu}=1e-5$. This provides a more interpretable view of the privacy-utility trade-off, where lower ($\epsilon_{\mu}$) values signify stronger, more practical privacy guarantees. The NVDP model consistently achieves better privacy-utility points than the VTDP ablation.
@@
This conclusion is strongly reinforced by the Rényi Divergence (RD) values in Table 1. As a direct measure of distinguishability, the RD reveals a substantial gap between NVDP and the baselines. In the same MRPC comparison, NVDP’s worst-case RD is 0.34. Conversely, VIB-fixed and VIB-learned exhibit much higher information leakage, with RD values of 2.98 and 2.14, respectively – nearly an order of magnitude higher. The benefit is particularly clear on SST-2; while NVDP achieves an RD of just 0.19, both VIB-fixed and VIB-learned exceed 2.20, and the VTDP ablation sits at 0.37. These RD results confirm that NVDP provides the tightest privacy guarantees for a given level of utility.
Together, these results demonstrate that NVDP is uniquely capable of preserving task-critical information while maintaining tight privacy guarantees. Its use of learned dynamic noise levels is much better than task-agnostic noise injection strategies at achieving good privacy guarantees while maintaining high utility. Its use of NVIB to manage the holistic noise level over an entire sequence shows consistent improvement over managing noise levels independently for each vector in a transformer’s multi-vector embeddings.
@@ 5 Conclusion
We propose a model that addresses the privacy concerns in sharing data for deep learning by integrating a nonparametric variational information bottleneck layer into the transformer architecture and sharing its transformer embeddings. By sharing embeddings sampled from the NVIB layer, our proposed nonparametric variational differential privacy model controls the amount of shared information despite there being many vectors in each transformer embedding, and this information is tailored to the needs of the target downstream task. We quantify the level of privacy with Rényi divergence differential privacy, and show that our NVDP model is able to share embeddings of data which provide strong privacy guarantees while maintaining good downstream model utility.
Our experimental results demonstrate the effectiveness of this approach by evaluating the privacy-utility tradeoff. We confirmed that NVDP consistently controls information leakage more effectively than a strong VIB-based ablation across a range of GLUE tasks. Crucially, by converting our Rényi Divergence measurements into interpretable $(\epsilon_{\mu},\lambda_{\mu})$-Bayesian Differential Privacy guarantees, we have shown that our model can achieve strong, practical privacy budgets while maintaining high model utility. This is a significant step towards deploying privacy-preserving transformer embeddings in real-world applications where clear and meaningful privacy assurances are required.
@@ Appendix A Additional Privacy Analysis
This appendix provides supplementary results to complement the analysis in the main text. Table 2 presents the comprehensive numerical results presented in Table 1, detailing the utility and privacy metrics for all tested KL-divergence regularization weights in Equation 5. So, this table provide the complete data used to generate the summary results in Table 1. Figure 2 visualize the privacy-utility trade-offs for the worst case max Rényi Divergence. These results demonstrate that our NVDP model consistently outperforms the VTDP ablation under all evaluation settings.
Table 2: Full experimental results for BERT-Base on GLUE, showing the complete privacy-utility trade-off across all tested KL-divergence regularization weights ($\lambda_{G},\lambda_{D}$). For each model, we report Utility (task-specific scores) and privacy metrics BDP ($\epsilon_{\mu}$) and RD. The RD metric are reported as max/avg, representing the worst-case and average-case divergence across all example pairs, respectively. Lower values are better for all privacy metrics ($\downarrow$).
@@
Figure 3: Accuracy versus maximum Rényi Divergence (RD) illustrating the worst-case privacy-utility tradeoff. Each point corresponds to a different KL regularization weight, where stronger regularization leads to better privacy (lower RD). The most favorable models are those closest to the upper-right corner.

### v2

@@ 4.2 Results on the GLUE Benchmark
Experimental Protocol
For each model, we perform five independent runs and select the best-performing run on the validation set for final evaluation on the test set. The NVDP and VIB models provide privacy by applying the same learned stochastic mapping at test time, ensuring that only noisy sanitized embeddings are shared. The utility of these private embeddings is measured by evaluating the classifier learned during training. The results in Table 1 represent the best privacy-utility trade-off found by sweeping across a range of regularization weights $\lambda_{G}$ and $\lambda_{D}$ (as defined in Equation 5). To quantify the privacy loss shown in Table 1, we fix the Rényi order to $\lambda=1.1$ (in Equation 7) and the BDP failure probability to $\delta_{\mu}=10^{-5}$ (in Equation 3), and report both the BDP and the maximum RD (RD max) across all test set pairs. Full results for all tested regularization strengths are provided in Appendix A.
Table 1: Privacy-utility trade-off on GLUE tasks for BERT-Base. We compare our proposed NVDP model against non-private baselines and private ablations (VIB-fixed, VIB-learned, and VTDP). For each private model, we report its best-achieved utility score alongside privacy guarantees. Privacy is measured via BDP (BDP $\downarrow$) and RD (RD $\downarrow$). Lower privacy values are better. The best-performing private model per task is bolded.
@@
Utility and Regularization Analysis
@@
We analyze this ability of our models to balance utility against privacy, to get a more complete understanding of the results from Table 1. Figure 2 plots accuracy against the BDP measure for the tunable models (NVDP and VTDP); detailed results are given in Table 2 in Appendix A. For this BDP guarantee, the results show a clear advantage for our proposed model. Across all datasets, the NVDP models (blue curves) consistently occupy the most favorable region of the plot – closest to the top-right corner – indicating a superior trade-off compared to the VTDP ablation (red curves). For example, on MRPC, to achieve a privacy budget comparable to NVDP (e.g., $\approx$ 10.6), the VTDP ablation’s accuracy drops to just 74.8%.
Figure 2: Accuracy versus BDP ($\epsilon_{\mu}$). The BDP budget ($\epsilon_{\mu}$) is calculated by finding the tightest privacy guarantee for a fixed $\delta_{\mu}=1e-5$. This provides a more interpretable view of the privacy-utility trade-off, where lower ($\epsilon_{\mu}$) values signify stronger, more practical privacy guarantees. The NVDP model consistently achieves better privacy-utility points than the VTDP ablation.
@@
This conclusion is strongly reinforced by the RD values in Table 1. As a direct measure of distinguishability, the RD reveals a substantial gap between NVDP and the baselines. In the same MRPC comparison, NVDP’s worst-case RD is 0.34. Conversely, VIB-fixed and VIB-learned exhibit much higher information leakage, with RD values of 2.98 and 2.14, respectively – nearly an order of magnitude higher. The benefit is particularly clear on SST-2; while NVDP achieves an RD of just 0.19, both VIB-fixed and VIB-learned exceed 2.20, and the VTDP ablation sits at 0.37. These RD results confirm that NVDP provides the tightest privacy guarantees for a given level of utility.
Together, these results demonstrate that NVDP is uniquely capable of preserving task-critical information while maintaining tight privacy guarantees. Its use of learned dynamic noise levels is much better than task-agnostic noise injection strategies at achieving good privacy guarantees while maintaining high utility. Its use of NVIB to manage the holistic noise level over an entire sequence shows consistent improvement over managing noise levels independently for each vector in a transformer’s multivector embeddings.
@@ 5 Conclusion
We propose a model that addresses the privacy concerns in sharing data for deep learning by integrating a NVIB layer into the transformer architecture and sharing its transformer embeddings. By sharing embeddings sampled from the NVIB layer, our proposed NVDP model controls the amount of shared information despite there being many vectors in each transformer embedding, and this information is tailored to the needs of the target downstream task. We quantify the level of privacy with RD DP, and show that our NVDP model is able to share embeddings of data which provide strong privacy guarantees while maintaining good downstream model utility.
Our experimental results demonstrate the effectiveness of this approach by evaluating the privacy-utility trade-off. We confirmed that NVDP consistently controls information leakage more effectively than a strong VIB-based ablation across a range of GLUE tasks. Crucially, by converting our RD measurements into interpretable $(\epsilon_{\mu},\lambda_{\mu})$-BDP guarantees, we have shown that our model can achieve strong, practical privacy budgets while maintaining high model utility. This is a significant step towards deploying privacy-preserving transformer embeddings in real-world applications where clear and meaningful privacy assurances are required.
@@ Appendix A Additional Privacy Analysis
This appendix provides supplementary results to complement the analysis in the main text. Table 2 presents the comprehensive numerical results presented in Table 1, detailing the utility and privacy metrics for all tested KL divergence regularization weights in Equation 5. So, this table provides the complete data used to generate the summary results in Table 1. Figure 2 visualizes the privacy-utility trade-offs for the worst-case max RD. These results demonstrate that our NVDP model consistently outperforms the VTDP ablation under all evaluation settings.
Table 2: Full experimental results for BERT-Base on GLUE, showing the complete privacy-utility trade-off across all tested KL divergence regularization weights ($\lambda_{G},\lambda_{D}$). For each model, we report Utility (task-specific scores) and privacy metrics BDP ($\epsilon_{\mu}$) and RD. The RD metric are reported as max/avg, representing the worst-case and average-case divergence across all example pairs, respectively. Lower values are better for all privacy metrics ($\downarrow$).
@@
Figure 3: Accuracy versus maximum RD illustrating the worst-case privacy-utility trade-off. Each point corresponds to a different KL regularization weight, where stronger regularization leads to better privacy (lower RD). The most favorable models are those closest to the upper-right corner.